OpenStudy (anonymous):

graph y= x2 + 1 using transformations, you would show the graph of y = x2 and the graph shifted up 1 unit. 푦= −2|x|+ 2

OpenStudy (compassionate):

나는 한국어를하지 않습니다

OpenStudy (compassionate):

Oh - you speak English!

OpenStudy (compassionate):

What exactly does, "푦," mean?

OpenStudy (anonymous):

Yes I do

OpenStudy (anonymous):

I have no idea, It was there when I uploaded my question

OpenStudy (anonymous):

you mean copied paste?

OpenStudy (compassionate):

Could you change it or go look at it again so we can understand your question?

OpenStudy (anonymous):

do you need me to redo the question I apologize for that

OpenStudy (compassionate):

Yes, please.

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

graph y= x2 + 1 using transformations, you would show the graph of y = x2 and the graph shifted up 1 unit. Please do not show only the final graph. 𝑦= −2|𝑥|+ 2

OpenStudy (anonymous):

I hope this is better for you to understand, thanks for the help!!

OpenStudy (compassionate):

I can't even see the characters.

OpenStudy (anonymous):

I post the question again in the reply box, do you see it?

OpenStudy (anonymous):

draw it instead using the draw button.

OpenStudy (whpalmer4):

I think the y=-2|x|+2 might be part of the next problem or something. y = x^2 is a parabola with vertex at the origin. y = x^2+1 is the same parabola, shifted up 1 unit, as seen here (dashed figure is original):